Joe has a mass of 50 kg and Sally has a mass of 30 kg. They sit in class with a separation of 0.5 meters. What is the gravitational attraction between Joe and Sally.
1
Expert's answer
2010-12-05T05:56:38-0500
According to the Newton's law of universal gravitation
F = GM1M2/R2
G = 6.674×10−11 N m2 kg−2
F = 6.674×10−11 * 50 * 30/ 0.52 = 4.0044x10-7 N